# Dual-VQE: A quantum algorithm to lower bound the ground-state energy

### v1

@@ Abstract
The variational quantum eigensolver (VQE) is a hybrid quantum–classical variational algorithm that produces an upper-bound estimate of the ground-state energy of a Hamiltonian. As quantum computers become more powerful and go beyond the reach of classical brute-force simulation, it is important to assess the quality of solutions produced by them. Here we propose a dual variational quantum eigensolver (dual-VQE) that produces a lower-bound estimate of the ground-state energy. As such, VQE and dual-VQE can serve as quality checks on their solutions; in the ideal case, the VQE upper bound and the dual-VQE lower bound form an interval containing the true optimal value of the ground-state energy. The idea behind dual-VQE is to employ semi-definite programming duality to rewrite the ground-state optimization problem as a constrained maximization problem, which itself can be bounded from below by an unconstrained optimization problem to be solved by a variational quantum algorithm. When using a convex combination ansatz in conjunction with a classical generative model, the quantum computational resources needed to evaluate the objective function of dual-VQE are no greater than those needed for that of VQE. We simulated the performance of dual-VQE on the transverse-field Ising model, and found that, for the example considered, while dual-VQE training is slower and noisier than VQE, it approaches the true value with error of order $10^{-2}$.
@@
Introduction—A key approach hoped to achieve quantum advantage on near-term quantum computers involves the use of variational quantum algorithms [5, 14]. The variational quantum eigensolver (VQE) was the first such algorithm proposed [36]: it estimates an upper bound on the ground-state energy of a Hamiltonian. The basic idea behind it is simple, based on the well known variational principle from physics [22]. For a Hamiltonian $H$, its ground-state energy is equal to its minimum eigenvalue $\lambda_{\min}(H)$, which is equal to the following minimization problem:
@@
VQE instead parameterizes the set of state vectors by a parameterized quantum circuit $U(\theta)$ acting on an easily preparable initial state $|\psi_{0}\rangle$, leading to $|\psi(\theta)\rangle\coloneqq U(\theta)|\psi_{0}\rangle$, and measures the expectation of each observable $H_{j}$ with respect to $|\psi(\theta)\rangle$ in order to obtain an estimate $\widehat{E_{j}}$ of $\langle\psi(\theta)|H_{j}|\psi(\theta)\rangle$. Then the sum $\sum_{j}\widehat{E_{j}}$ is an estimate of the desired expectation $\langle\psi(\theta)|H|\psi(\theta)\rangle$. The estimates of $\langle\psi(\theta)|H|\psi(\theta)\rangle$ and its gradient can be fed to a classical optimization procedure and this process iterated until convergence or a maximum number of iterations is reached. As a consequence of (3), this procedure in the ideal case produces an estimate of an upper bound for the ground-state energy $\lambda_{\min}(H)$.
@@
Although there have been various classical algorithms proposed for obtaining lower bounds on the ground-state energy [28, 29, 30, 8, 23], they all ultimately suffer from the exponential increase of the space on which $H$ acts, as the number of particles increases. Our approach is fundamentally different, involving a reformulation of the original optimization problem in terms of its SDP dual, and then bounding that quantity from below using a variational quantum algorithm. Our approach is also complementary to [9, Section VI-A], which does not take a variational approach, nor does it provably obtain a lower bound on the ground-state energy.
@@
The rest of our paper proceeds by detailing the dual-VQE method outlined above, and then we report the results of simulations of the transverse-field Ising model, a key example of physical interest. We find that the dual-VQE approach on a two-qubit example takes more iterations to converge as compared to the standard VQE optimization; however, it approaches the true value with error on the order of $10^{-2}$. We also note that in this example, the objective function values across training are noisier for dual-VQE when compared to VQE. We expect this behavior due to the increase in number of terms and parameters when estimating the ground-state energy with dual-VQE. Details of numerical simulations are given below, and training plots are shown in Figure 1.
@@
Dual-VQE theoretical derivation—Let us first use some basic SDP duality theory to derive a well known result, that the following weak-duality inequality holds:
$\sup_{\eta\in\mathbb{R}}\left\{\eta:\eta I\leq H\right\}\leq\inf_{\rho\in\mathcal{D}}\operatorname{Tr}[H\rho],$ (4)
@@
where $\sup_{\eta\in\mathbb{R}}\left\{\eta:\eta I\leq H\right\}$ is the SDP dual of (1). Note that actually equality holds in (4) due to Slater’s theorem [13]: we can pick $\rho=I/d$, where $d$ is the dimension of the underlying Hilbert space, and $\eta<\lambda_{\min}(H)$ as strictly feasible choices in both the dual and primal, thus satisfying the conditions of Slater’s theorem. Recall that the matrix inequality $A\geq B$ holds for Hermitian matrices $A$ and $B$ if $A-B$ is a positive semi-definite matrix. Then consider that
@@
The second equality follows by introducing the Lagrange multiplier $\eta$ and because the constraint $\operatorname{Tr}[\rho]=1$ does not hold if and only if $\sup_{\eta\in\mathbb{R}}\left\{\eta\left(1-\operatorname{Tr}[\rho]\right)\right\}=+\infty$. The third equality holds by simple algebra. The sole inequality follows from the standard max-min inequality. The final equality follows by thinking of $\rho$ as a matrix Lagrange multiplier and because the constraint $\eta I\leq H$ does not hold if and only if $\inf_{\rho\geq 0}\left\{\operatorname{Tr}[\left(H-\eta I\right)\rho]\right\}=-\infty$.
Now that we have recalled this standard result, let us focus on showing how to estimate the lower bound in (4) on a quantum computer. Recalling that the matrix inequality $\eta I\leq H$ is equivalent to the existence of a positive semi-definite matrix $W$ such that $H-\eta I=W$, we conclude that
@@
Our next observation is that every positive semi-definite matrix $W\neq 0$ can be written as a scaled density matrix, i.e., $W=\nu\omega$, where $\nu=\operatorname{Tr}[W]$ and $\omega=W/\nu\in\mathcal{D}$, implying that
@@
where $c>0$ is a penalty parameter and $\left\|Z\right\|_{2}\coloneqq\sqrt{\operatorname{Tr}[Z^{{\dagger}}Z]}$ is the Hilbert–Schmidt norm of a matrix $Z$. Note that the equality in (12) holds whenever the penalty term is faithful (specifically, we invoke [6, Proposition 5.2.1] to justify this conclusion, setting $\lambda_{k}=0$ therein for all $k$). By faithfulness, we mean that the penalty term is equal to zero if and only if $H-\eta I=\nu\omega$. This is true for the Hilbert–Schmidt norm, and this particular norm has the convenient advantage that it can be estimated by quantum sampling procedures that we recall below.
@@
Piecing together all these steps in (4)–(13), expanding the Hilbert-Schmidt norm, and defining
@@
[…]
Dual-VQE algorithm for lower bounding the ground-state energy—The basic idea behind the dual-VQE algorithm is to follow the standard penalty method [6, Section 5.2]: define a sequence $\left(f^{\ast}(c_{k})\right)_{k\in\mathbb{N}}$ of optimization problems where $\left(c_{k}\right)_{k\in\mathbb{N}}$ is a monotone increasing sequence of penalty parameters. We then solve the optimization for each $k$ and use the solution as the initial guess for the next iteration. Under the assumption that each optimization $f^{\ast}(c_{k})$ can be solved, the method is guaranteed to converge to the correct solution. In our simulations, we used a constant value of the penalty parameter $c$ and found that it performs well in practice.
In dual-VQE, we estimate the expectation $\operatorname{Tr}[H\omega(\theta)]$ by sampling from a quantum computer, and we use other approaches, either quantum or classical, for estimating the purity $\operatorname{Tr}[\left(\omega(\theta)\right)^{2}]$. If employing gradient descent for solving each optimization, one should estimate the gradient as well, which we discuss further below. The main additional overhead compared to VQE is the need to employ an ansatz for generating a mixed state $\omega(\theta)$ and to estimate the purity $\operatorname{Tr}[\left(\omega(\theta)\right)^{2}]$, given that VQE already estimates a term like $\operatorname{Tr}[H\omega(\theta)]$. However, as we discuss below, one approach reduces the estimation of the purity $\operatorname{Tr}[\left(\omega(\theta)\right)^{2}]$ to a classical sampling task, thus removing the need for a quantum computer for this part.
There are at least two general ways of parameterizing the set of mixed states, called the purification ansatz [16, 35, 19] and the convex combination ansatz [46, 26, 19, 41]. Focusing on the first one to start, it is based on the purification principle [12, 44, 45], in which every mixed state is realized as the marginal of a pure state on a larger system. For the purification ansatz, we simply start with a parameterized family $\left\{|\psi(\theta)\rangle_{RS}\right\}_{\theta\in\Theta}$ of pure states of a reference system $R$ and the system $S$ of interest, and then we obtain a parameterized family $\left\{\omega(\theta)\right\}_{\theta\in\Theta}$ of mixed states via the partial trace $\omega(\theta)=\operatorname{Tr}_{R}[|\psi(\theta)\rangle\!\langle\psi(\theta)|_{RS}]$. We can estimate the purity $\operatorname{Tr}[\left(\omega(\theta)\right)^{2}]$ by means of the destructive swap test [21] (see, e.g., [11, Section 2.2] for a precise statement of the algorithm). We can estimate the gradient of the objective function in (16) by means of the parameter-shift rule [27, 33, 39], simply because
@@
$F$ is the unitary and Hermitian swap operator, and we have suppressed system labels in (22) for brevity (see also [19, Eq. (C2)] in this context).
@@
Interestingly, although it is necessary to use a quantum computer to estimate the purity $\operatorname{Tr}[\left(\omega(\theta)\right)^{2}]$ for the purification ansatz, it is not necessary to do so when using the convex combination ansatz in conjunction with a neural-network-based generative model: it suffices to employ a classical sampling approach, thus further reducing the quantum resource requirements for dual-VQE (see also [19, Eqs. (20)–(23)] in this context). This is because $\operatorname{Tr}[\left(\omega(\varphi,\gamma)\right)^{2}]=\sum_{x}p_{\varphi}^{2}(x)$, so that one can repeatedly sample from $p_{\varphi}$ and perform a collision test to estimate $\sum_{x}p_{\varphi}^{2}(x)$. Thus, for the convex combination ansatz used in the aforementioned way, one only needs a quantum computer to estimate the expectation
@@
[…]

### v2

@@ Abstract
The variational quantum eigensolver (VQE) is a hybrid quantum–classical variational algorithm that produces an upper-bound estimate of the ground-state energy of a Hamiltonian. As quantum computers become more powerful and go beyond the reach of classical brute-force simulation, it is important to assess the quality of solutions produced by them. Here we propose a dual variational quantum eigensolver (dual-VQE) that produces a lower-bound estimate of the ground-state energy. As such, VQE and dual-VQE can serve as quality checks on their solutions; in the ideal case, the VQE upper bound and the dual-VQE lower bound form an interval containing the true optimal value of the ground-state energy. The idea behind dual-VQE is to employ semi-definite programming duality to rewrite the ground-state optimization problem as a constrained maximization problem, which itself can be bounded from below by an unconstrained optimization problem to be solved by a variational quantum algorithm. When using a convex combination ansatz in conjunction with a classical generative model, the quantum computational resources needed to evaluate the objective function of dual-VQE are no greater than those needed for that of VQE. We also show that the problem is well suited for classical pretraining using matrix product states and these methods help warm-start the optimization. We simulated the performance of dual-VQE on the transverse-field Ising model with and without pretraining and found that, for the example considered, while dual-VQE training is slower and noisier than VQE, it approaches the true value with an error of order $10^{-2}$.
@@
I Introduction
I.1 Background
A key approach hoped to achieve quantum advantage on near-term quantum computers involves the use of variational quantum algorithms [1, 2]. The variational quantum eigensolver (VQE) was the first such algorithm proposed [3]: it estimates an upper bound on the ground-state energy of a Hamiltonian. The basic idea behind it is simple, based on the well known variational principle from physics [4]. For a Hamiltonian $H$, its ground-state energy is equal to its minimum eigenvalue $\lambda_{\min}(H)$, which is equal to the following minimization problem:
@@ I.1 Background
VQE instead parameterizes the set of state vectors by a parameterized quantum circuit $U(\theta)$ acting on an easily preparable initial state $|\psi_{0}\rangle$, leading to $|\psi(\theta)\rangle\coloneqq U(\theta)|\psi_{0}\rangle$, and measures the expectation of each observable $H_{j}$ with respect to $|\psi(\theta)\rangle$ in order to obtain an estimate $\widehat{E_{j}}$ of $\langle\psi(\theta)|H_{j}|\psi(\theta)\rangle$. Then the sum $\sum_{j}\widehat{E_{j}}$ is an estimate of the desired expectation $\langle\psi(\theta)|H|\psi(\theta)\rangle$. The estimates of $\langle\psi(\theta)|H|\psi(\theta)\rangle$ and its gradient can be fed to a classical optimization procedure, and this process iterated until convergence or a maximum number of iterations is reached. As a consequence of (3), this procedure in the ideal case produces an estimate of an upper bound on the ground-state energy $\lambda_{\min}(H)$.
I.2 Main contribution: Dual-VQE method
@@
Although there have been various classical algorithms proposed for obtaining lower bounds on the ground-state energy [9, 10, 11, 12, 13], they all ultimately suffer from the exponential increase of the space on which $H$ acts, as the number of particles increases. Our approach is fundamentally different, involving a reformulation of the original optimization problem in terms of its SDP dual, and then bounding that quantity from below using a variational quantum algorithm. Our approach is also complementary to [14, Section VI-A], which does not take a variational approach, nor does it provably obtain a lower bound on the ground-state energy.
I.3 Pretraining with matrix product states
A key issue that plagues variational quantum algorithms is the presence of barren plateaus in the training landscape [15, 16, 17], meaning that the gradient (or equivalently cost differences [18]) vanishes exponentially fast in the number of qubits. At their core, barren plateaus arise from a curse of dimensionality [19] stemming from the exponentially scaling size of the Hilbert space. On the other hand, standard techniques for provably avoiding barren plateaus do so by encoding the relevant dynamics in a smaller subspace making it effectively possible to classically simulate the parameterised quantum circuit [19].
An important question for dual-VQE, shared also with other variational quantum algorithms, is thus whether we can identify quantum models that both avoid barren plateaus and cannot be classically simulated. A promising avenue to do so is to note that theoretical analysis of barren plateaus and the simulability of variational quantum algorithms is confined to average case analyses. It is possible that there might exist clever initialization strategies to explore small subsections of the landscape that exhibit substantial gradients, and are also hard to simulate classically.
Motivated as such, in this work we develop a matrix product state (MPS) pretrainer that uses tensor networks to initialize the ansatz parameters in a more efficient way [20, 21]. We find that this MPS pretrainer allows us to start the quantum optimization from a favorable initial point, compared to random initializations, as evidenced by the convergence of the cost function value closer to the true value (in this context, see Figure 3).
@@ I.3 Pretraining with matrix product states
We find that, without pretraining, the dual-VQE approach on a two-qubit example takes more iterations to converge as compared to the standard VQE optimization; however, it approaches the true value with error on the order of $10^{-2}$. We also note that in this example, the objective function values across training are noisier for dual-VQE when compared to VQE. We expect this behavior due to the increase in number of terms and parameters when estimating the ground-state energy with dual-VQE. Details of numerical simulations are given below, and training plots are shown in Figure 2.
With the introduction of pretraining, we tackle a larger problem and find that, as expected, pretraining with a larger bond dimension allows for more efficient optimization. Concretely, with a maximum bond dimension of eight, we find the ground state of a three-qubit model with a relative error of $0.5\%$.
I.4 Paper organization
The rest of our paper proceeds by detailing the dual-VQE method outlined above. In Section II, we first describe the theoretical derivation of the key result. We then follow this up in Section III with more intricate details of the simulations, including the penalty value, which plays a key role in the theoretical derivation, the ansatz choices, and the pretraining method that we employ. Finally, in Section IV, we report the results of simulations of the transverse-field Ising model, a key example of physical interest. Appendix A provides background on matrix product states, and Appendix B provides further details of the pretraining method that we used.
II Theoretical derivation
@@
Let us first apply basic SDP duality theory to derive a well known result, that the following strong-duality equality holds for the optimization problem in (1):
$\sup_{\eta\in\mathbb{R}}\left\{\eta:\eta I\leq H\right\}=\inf_{\rho\in\mathcal{D}}\operatorname{Tr}[H\rho],$ (4)
@@
where $\sup_{\eta\in\mathbb{R}}\left\{\eta:\eta I\leq H\right\}$ is the SDP dual of (1). To see this, recall that the matrix inequality $A\geq B$ holds for Hermitian matrices $A$ and $B$ if ABA-B is a positive semi-definite matrix. Then consider that
@@
The second equality follows by introducing the Lagrange multiplier $\eta$ and because the constraint $\operatorname{Tr}[\rho]=1$ does not hold if and only if $\sup_{\eta\in\mathbb{R}}\left\{\eta\left(1-\operatorname{Tr}[\rho]\right)\right\}=+\infty$. The third equality holds by simple algebra. The sole inequality follows from the standard max-min inequality. The final equality follows by thinking of $\rho$ as a matrix Lagrange multiplier and because the constraint $\eta I\leq H$ does not hold if and only if $\inf_{\rho\geq 0}\left\{\operatorname{Tr}[\left(H-\eta I\right)\rho]\right\}=$-$\infty$. Thus, we have established the weak-duality inequality
$\sup_{\eta\in\mathbb{R}}\left\{\eta:\eta I\leq H\right\}\leq\inf_{\rho\in\mathcal{D}}\operatorname{Tr}[H\rho].$ (10)
Note that equality holds in (4) due to Slater’s theorem [22]: we can pick $\rho=I/d$, where $d$ is the dimension of the underlying Hilbert space, and $\eta<\lambda_{\min}(H)$ as strictly feasible choices in both the dual and primal, thus satisfying the conditions of Slater’s theorem.
Now that we have recalled this standard result, let us focus on showing how to estimate the left-hand side of (4) on a quantum computer. Recalling that the matrix inequality $\eta I\leq H$ is equivalent to the existence of a positive semi-definite matrix $W$ such that $H-\eta I=W$, we conclude that
@@
Our next observation, related to observations used in [23], is that every positive semi-definite matrix $W\neq 0$ can be written as a scaled density matrix, i.e., $W=\nu\omega$, where $\nu=\operatorname{Tr}[W]$ and $\omega=W/\nu\in\mathcal{D}$, implying that
@@
where $c>0$ is a penalty parameter and $\left\|Z\right\|_{2}\coloneqq\sqrt{\operatorname{Tr}[Z^{{\dagger}}Z]}$ is the Hilbert–Schmidt norm of a matrix $Z$. Note that the equality in (13) holds whenever the penalty term is faithful (specifically, we invoke [24, Proposition 5.2.1] to justify this conclusion, setting $\lambda_{k}=0$ therein for all $k$). By faithfulness, we mean that the penalty term is equal to zero if and only if $H-\eta I=\nu\omega$. This is true for the Hilbert–Schmidt norm, and this particular norm has the convenient advantage that it can be estimated by quantum sampling procedures that we recall below.
@@
Piecing together all these steps in (4)–(14), expanding the Hilbert–Schmidt norm, and defining
@@
In the rest of the work, we use the term “objective function” to refer to $\eta$, and “penalty” to refer to the sum of all terms in (16) that are multiplied by the penalty parameter $c$, i.e.,
$\operatorname{Tr}[H^{2}]+\eta^{2}2^{n}+\nu^{2}\operatorname{Tr}[\omega(\theta)^{2}]\\ -2\eta\operatorname{Tr}[H]-2\nu\operatorname{Tr}[H\omega(\theta)]+2\eta\nu.$ (18)
III Algorithm construction
The basic idea behind the dual-VQE algorithm is to follow the standard penalty method approximately [24, Section 5.2]: define a sequence $\left(f^{\ast}(c_{k})\right)_{k\in\mathbb{N}}$ of optimization problems where $\left(c_{k}\right)_{k\in\mathbb{N}}$ is a monotone increasing sequence of penalty parameters. We then solve the optimization for each $k$ and use the solution as the initial guess for the next iteration. Under the assumption that each optimization $f^{\ast}(c_{k})$ can be solved, the method is guaranteed to converge to the correct solution. In our simulations, we use a variety of update methods to change the value of $c$ as the optimization progresses. In some cases, we pick $c$ to be a high value and keep it constant throughout the optimization. In other cases, we begin with a smaller value, and ramp it up as the iteration count grows.
@@ III Algorithm construction
As is common in Hamiltonian models relevant for physics and chemistry, the number of non-zero coefficients in the tuple $\left(\alpha_{\overrightarrow{x}}\right)_{\overrightarrow{x}}$ is polynomial inn~n, so that the objective function in (16) can be efficiently estimated. In this work, we focus on the transverse-field Ising model given in (32), which falls under this category.
In dual-VQE, we estimate the expectation $\operatorname{Tr}[H\omega(\theta)]$ by sampling from a quantum computer, and we use other approaches, either quantum or classical, for estimating the purity $P\equiv\operatorname{Tr}\!\left[\left(\omega(\theta)\right)^{2}\right]$. If employing gradient descent for solving each optimization, one should estimate the gradient as well, which we discuss further below. The main additional overhead compared to VQE is the need to employ an ansatz for generating a mixed state $\omega(\theta)$ and to estimate the purity $P$, given that VQE already estimates a term like $\operatorname{Tr}[H\omega(\theta)]$. However, as we discuss below, one approach reduces the estimation of the purity $P$ to a classical sampling task, thus removing the need for a quantum computer for this part.
III.1 Parameterization
There are at least two general ways of parameterizing the set of mixed states, called the purification ansatz [25, 23, 26] and the convex combination ansatz [27, 28, 26, 29] (see also [8]). Focusing on the first one to start, it is based on the purification principle [30, 31, 32], in which every mixed state is realized as the marginal of a pure state on a larger system. For the purification ansatz, we simply start with a parameterized family $\left\{|\psi(\theta)\rangle_{RS}\right\}_{\theta\in\Theta}$ of pure states of a reference system $R$ and the system $S$ of interest, and then we obtain a parameterized family $\left\{\omega(\theta)\right\}_{\theta\in\Theta}$ of mixed states via the partial trace $\omega(\theta)=\operatorname{Tr}_{R}[|\psi(\theta)\rangle\!\langle\psi(\theta)|_{RS}]$. We can estimate the purity $\operatorname{Tr}[\left(\omega(\theta)\right)^{2}]$ by means of the destructive swap test [7] (see, e.g., [33, Section 2.2] for a precise statement of the algorithm). We can estimate the gradient of the objective function in (17) by means of the parameter-shift rule [34, 35, 36], simply because
@@ III.1 Parameterization
$F$ is the unitary and Hermitian swap operator, and we have suppressed some system labels in (24) for brevity (see also [26, Eq. (C2)] in this context).
@@
Interestingly, although it is necessary to use a quantum computer to estimate the purity $\operatorname{Tr}[\left(\omega(\theta)\right)^{2}]$ for the purification ansatz, it is not necessary to do so when using the convex combination ansatz in conjunction with a neural-network-based generative model: it suffices to employ a classical sampling approach, thus further reducing the quantum resource requirements for dual-VQE (see also [26, Eqs. (20)–(23)] in this context). This is because $\operatorname{Tr}[\left(\omega(\varphi,\gamma)\right)^{2}]=\sum_{x}p_{\varphi}^{2}(x)$, so that one can repeatedly sample from $p_{\varphi}$ and perform a collision test to estimate $\sum_{x}p_{\varphi}^{2}(x)$. Thus, for the convex combination ansatz used in the aforementioned way, one only needs a quantum computer to estimate the expectation
@@
which can be done by repeating the following procedure and calculating the sample mean: take a samplex~x from $p_{\varphi}(x)$, prepare the state $U(\gamma)|x\rangle$, measure this state according to $H$, and record the outcome. To estimate the gradient of the purity $\sum_{x}p_{\varphi}^{2}(x)$ and the gradient of the expectation $\operatorname{Tr}[H\omega(\varphi,\gamma)]$ with respect to $\varphi$, perhaps the simplest approach is to use the simultaneous perturbation stochastic approximation [41], because the distribution $p_{\varphi}(x)$ in this case is not differentiable, and as such, standard methods like backpropagation are not readily applicable (see [38] for further discussions and other approaches).
III.2 Pretraining
Pretraining broadly refers to a set of relatively computational inexpensive techniques that kick start (or warm start) an optimization process, as opposed to a random start. In several domains, pretraining has been shown to be advantageous in reducing the training time and accelerating convergence [42, 43].
In this section, we describe the details of the particular pretraining process used in this work. The overall process consists of two phases: pretraining, followed by a translation into a quantum circuit. The pretraining is conducted using matrix product states (MPS) [44], and the final translation consists of a combination of two techniques – analytic decomposition (AD) and optimizing decomposition (OD), as proposed in [45, 46]. Important preliminaries on tensor networks and matrix product states can be found in Appendix A, and key ideas of the pretraining process, including an in-depth review of AD and OD, can be found in Appendix B.
Before delving into the specifics of the algorithm, we aim to establish the rationale behind pretraining using MPS. Our motivation for this approach stems from Refs. [47, 21], which explored solving the Variational Quantum Eigensolver (VQE) problem by pretraining it with MPS. The study demonstrated that parameterized quantum circuits for VQE exhibited significant performance improvements when used in conjunction with a pretrained initialization, in contrast to a random initialization. When applied to the dual-VQE problem, we find that pretraining provides a significant kick start, with the advantage increasing with maximum bond dimension of the MPS.
The overall process consists of three phases: pretraining, translation, and quantum training. Since it is more complex to map an MPS to a convex combination ansatz state, we used the purification ansatz in our pretraining procedure. We now delve into the three stages in more detail.
III.2.1 Pretraining using matrix product states
The dual-VQE problem (without the final parameterization of the state $\omega$) is given by
$\sup_{\begin{subarray}{c}\eta\in\mathbb{R},\nu\geq 0,\\ \omega\in\mathcal{D}\end{subarray}}f(\eta,\nu,\omega,c),$ (28)
where the function $f$ is the dual-VQE cost function and is given by
$f(\eta,\nu,\omega,c)\coloneqq\eta-c\big(\operatorname{Tr}[H^{2}]+\eta^{2}2^{n}+\nu^{2}\operatorname{Tr}[\omega^{2}]\\ -2\eta\operatorname{Tr}[H]-2\nu\operatorname{Tr}[H\omega]+2\eta\nu\big).$ (29)
Figure 1: (a) Realizing mixed states using MPSs. The red tensors are the reference qubits, and the blue tensors are the system qubits. (b) Tensor network for estimating $\operatorname{Tr}[H\omega]$. The tensors in yellow represent the Pauli strings in the decomposition of $H$. (c) Tensor network for estimating $\operatorname{Tr}[\omega^{2}]$.
As discussed in Appendix A, the mixed state $\omega$ is realized as the partial trace of an MPS on a larger system. The MPS in this case represents the purification of the mixed state of interest. A schematic of this construction is given in Figure 1a).
We begin with a randomly initialized MPS with $n_{S}+n_{R}$ nodes in the canonical form, where $n_{R}$ is the number of qubits in the reference system $R$, and $n_{S}$ is the number of qubits in the system $S$ of interest. Again, we pick $n_{S}$ to be at least as large as $n_{R}$ to ensure a full-rank mixed state on the system qubits. The two terms that depend on $\omega$ are $\operatorname{Tr}[H\omega]$ and $\operatorname{Tr}[\omega^{2}]$. These terms are estimated using the tensor networks in Figure 1b) and Figure 1c), respectively.
In our work, we use the quimb package to realize matrix product states [48]. The package allows us to randomly initialize MPSs, contract specific indices, and compute gradients using automatic differentiation.
We begin all pretraining with a randomly initialized MPS where each entry of all the tensors are drawn from the normal distribution with zero mean and unit variance, and we use gradient ascent to train the parameters of the MPS to maximize the cost function in (28).
III.2.2 Convert final matrix product state to parameterized quantum circuit
After pretraining, we convert the resulting MPS into a parameterized quantum circuit composed of one- and two-qubit gates. If the MPS has a maximum bond dimension of two ($\chi=2$), it is exactly decomposable into a single layer of two-qubit unitaries [45]. However, for an arbitrary MPS with $\chi>2$, an exact decomposition into two-qubit gates is not generally feasible. This is because each MPS bond with dimension $\chi>2$ corresponds to a unitary acting on $\lceil\log_{2}(\chi)\rceil+1$ qubits, and it is not always possible to decompose any multi-qubit unitary exactly into a sequence of two-qubit gates.
To translate the final MPS into a parameterized quantum circuit, we make use of the algorithms put forth in [46]. The building blocks of the different algorithms are called the Analytical Decomposition (AD) algorithm and Optimizing Decomposition (OD). We note that in this work and in prior works, the AD algorithm is also called the disentangling algorithm.
For the sake of clarity and completeness, we provide a detailed description of both the AD and OD algorithms in Appendix B.
Remark 2
In [46], the OD algorithm consists of multiple iterations of training, where a unitary $U$ is replaced with a better unitary $U^{\prime}$. To find this unitary $U^{\prime}$, the algorithm uses the singular value decomposition (SVD) to find an optimal unitary $U_{\operatorname{new}}$, but it uses a small learning rate to prevent large updates. The update rule to replace unitary $U$ is as follows:
$U^{\prime}=U(U^{\dagger}U_{\operatorname{new}})^{\beta},$ (30)
where $\beta\in(0,1)$ is the learning rate, appropriately chosen beforehand. In this work, we introduce a different update rule, given by
$U^{\prime}=(1-\beta)\,U+\beta\,U_{\operatorname{new}}.$ (31)
We note here that, as defined above, $U^{\prime}$ is not unitary. To eliminate this issue, we perform an SVD of $U^{\prime}$, obtaining $U^{\prime}=WDV^{{\dagger}}$ and we then set $U^{\prime}\leftarrow WV^{{\dagger}}$. In our numerical investigations, we observed that the update rule in (31) leads to faster convergence and better overall fidelity compared to using (30).
In [46], the authors combined both the OD and AD algorithms in six different ways to translate the final trained MPS into a quantum circuit. From their work, we use three of the algorithms that they refer to as $D_{\operatorname{all}}$, $O_{\operatorname{all}}$, and $D_{\operatorname{all}}O_{\operatorname{all}}$.
  • $D_{\operatorname{all}}$: Consists of the AD algorithm only. Multiple layers are created using the AD algorithm with no optimization.
  • $O_{\operatorname{all}}$: Consists of the OD algorithm only. Multiple layers are randomly initalized and then optimized using the OD algorithm.
  • $D_{\operatorname{all}}O_{\operatorname{all}}$: Begins with the AD algorithm as an initial circuit and then uses the OD algorithm to further improve fidelity.
We run all three variants above and finally pick the translated circuit with the highest fidelity.
While the translated circuit is now composed of only two-qubit unitaries, it still needs to be cast into a form realizable on a quantum device. To that end, we utilize the KAK decomposition to decompose each two-qubit unitary into a product of parameterized rotation gates [49]. Each two-qubit unitary contributes a total of 15 parameters.
III.2.3 Train parameterized quantum circuit until convergence
Lastly, we use the translated MPS to warm-start the quantum optimization. Before we begin the quantum training, we append additional layers of the parameterized ansatz that are trivially initialized (identity operations). Then, all the parameters, including those from the pretrained portion of the circuit, are trained.
IV Results
We now present the various results of simulations that test the performance of dual-VQE on the transverse-field, one-dimensional Ising model, described by the following Hamiltonian of $n$ qubits:
$H_{\operatorname{TFI}}\coloneqq\sum_{i=1}^{n-1}J_{i}\sigma_{Z}^{i}\otimes\sigma_{Z}^{i+1}+\sum_{i=1}^{n}g_{i}\sigma_{X}^{i},$ (32)
where tensor products with identities are left implicit, $J_{i}$ is an interaction strength, and $g_{i}$ is related to the strength of the external field. In particular, we showcase our algorithm on the following example problem instance:
$H=\sum_{i=1}^{n-1}\sigma_{Z}^{i}\otimes\sigma_{Z}^{i+1}+\sum_{i=1}^{n}\sigma_{X}^{i},$ (33)
for $n=2$ and $n=3$ qubits. We simulated the dual-VQE problem using three different ansätze – the purification ansatz, the convex combination ansatz, and finally, MPS pretraining combined with a purification ansatz. We now present the details of the individual experiments and corresponding plots.